OpenStudy (australopithecus):

A Styrofoam slab has thickness h and density ρs. When a swimmer of mass m is resting on it, the slab floats in fresh water with its top at the same level as the water surface. Find the area of the slab. I assume the force exerted by the water needs to be equal to the force exerted by gravity on the object? so I have where m is the mas of the swimmer M is the mass of the Styrofoam (M + m)g = B since M = ρA we have that (ρA + m)g = B then (ρA + m) = B/g = A = (B/g - m)/ρ I'm pretty sure I'm wrong anyone able to help?

OpenStudy (australopithecus):

I mean I cant use this method to solve this because I dont know what B is

OpenStudy (australopithecus):

so I can just use A = ((m+ M)g/g - m)/ρ A = (M)/p

OpenStudy (australopithecus):

probably wrong

OpenStudy (shane_b):

Well, this part is obviously right: (M + m)g = B I'm still thinking about the rest

OpenStudy (shane_b):

Maybe I'm still sleepy but how did you figure that M = pA

OpenStudy (australopithecus):

that is just a general formula

OpenStudy (australopithecus):

density = mass/volume densityV = mass V = A

OpenStudy (shane_b):

Nm....I was thinking pressure not density. Yep, got it.

OpenStudy (shane_b):

I really don't see anything wrong here

OpenStudy (australopithecus):

yeah but I assume the height comes into it

OpenStudy (shane_b):

I think you're right...I need paper. One min

OpenStudy (australopithecus):

thanks I appreciate the help

OpenStudy (shane_b):

No help just yet :P

OpenStudy (shane_b):

I think I have it...it will take a minute to write this out though.

OpenStudy (australopithecus):

k thanks

OpenStudy (shane_b):

Starting:\[(M+m)g=B\]\[M+m=\frac{B}{g}\]\[ M=\rho V=\rho Ah\]\[\rho Ah+m=\frac{B}{g}\]\[\rho Ah=\frac{B}{g}-m\]\[A=\frac{\frac{B}{g}-m}{\rho h}\]Check it but I think that's correct.

OpenStudy (shane_b):

I used the normal density formula...not surface density.

OpenStudy (australopithecus):

give me a second getting error

OpenStudy (australopithecus):

why is pV = pAh ?

OpenStudy (shane_b):

That's where the mistake was made...I'm aggravated at myself for not seeing it immediately. V = area x height

OpenStudy (australopithecus):

oh true

OpenStudy (shane_b):

You had it right initially...except for that part.

OpenStudy (australopithecus):

thanks :)

OpenStudy (shane_b):

you're welcome :)